OpenStudy (anonymous):

2sec x + = 1 , Solve this on the interval [0,2Pi) Please answer and explain

OpenStudy (anonymous):

2sec x + ? =1 ??

OpenStudy (anonymous):

2sec x + 5 = 1 ...sorry

OpenStudy (anonymous):

@dyare

OpenStudy (anonymous):

2sec x=1-5 2sec x=-4 sec x=-2 x=inverse sec(-2) x= 1/inv. cos(-2)

OpenStudy (anonymous):

Available answers are A. 7PI/6, 11PI/6 B. 2PI/3, 4PI/3 C. Pi/3, 5PI/3 D. Pi/6, 5pi /6

OpenStudy (anonymous):

@ajprincess :3

OpenStudy (shubhamsrg):

secx = -2 cos x = -1/2 x= arccos(-1/2) x= 2pi/3 and 4pi/3

OpenStudy (anonymous):

tysm

OpenStudy (shubhamsrg):

tysm ?

OpenStudy (shubhamsrg):

thank you super man ?? o.O :P